Anne is 16 years younger than Max. in 5 years, Max will be twice as old as Anne will be then. Find their ages now.
Found 2 solutions by CubeyThePenguin, josgarithmetic:
Answer by CubeyThePenguin(3113)

a = Anne's age
m = Max's age

a = m - 16
m + 5 = 2(a + 5)

Substitute the first equation into the second.

m + 5 = 2(m - 16 + 5)
m + 5 = 2(m - 11)
m + 5 = 2m - 22
27 = m

Max is 27 and Anne is 11.

Answer by josgarithmetic(39617)